tan(alfa+beta)=1/2,tan(alfa-bata)=1/3, then find the value of tan2alfa

Grade:12th pass

1 Answers

poojitha
6 years ago
tan[(alfa+beta)+(alfa-beta)] = [tan(alfa+beta)+tan(alfa-beta)]/[1-tan(alfa+beta).tan(alfa-beta)]                       =(1/2+1/3)/(1-1/2.1/3) = (5/6)/(5/6) = 1         [tan2alfa=tan(alfa+beta+alfa-beta)]